Your family likes to go to baseball games. At one game your family bought 5 soft drinks and 5 hot dogs for $22.25. At the next game your family attended they bought 4 soft drinks and 3 hot dogs for $14.50. What is the cost of one soft drink and one hot dog?
Let s = the cost of one soft drink
Let h = the cost of one hot dog
a. Write a system of equations modeling the situation described above.
b. Solve the system for the cost of one soft drink + one hot dog. (You will have to add your solutions together for this one.)

Answer by stanbon(29602)

family bought 5 soft drinks and 5 hot dogs for $22.25. At the next game your family attended they bought 4 soft drinks and 3 hot dogs for $14.50. What is the cost of one soft drink and one hot dog?
Let s = the cost of one soft drink
Let h = the cost of one hot dog
a. Write a system of equations modeling the situation described above.
1st: 5s + 5h = 22.25
2nd: 4s + 3h = 14.50
-----------------------------
b. Solve the system for the cost of one soft drink + one hot dog.
Multiply 1st by 3 and 2nd by 5 to get:
3rd: 15s + 15h = 66.75
4th: 20s + 15h = 72.50
--------------
Subtract 3rd from 4th to get:
5s = 5.75
s = 1.15 (price of a solf drink)
---------------------
Substitute into 1st to solve for "h".
5(1.15)+5h = 22.25
5.75 +5h = 22.25
5h = 16.50
h = 3.50 (price of a hot dog)
==========================
Cheers,
Stan H.